# Meta-Learning Neural Process for Implied Volatility Surfaces with SABR-induced Priors

### v1

@@ Abstract
Constructing the implied volatility surface (IVS) is reframed as a meta-learning problem—training across trading days to learn a general process that reconstructs a full IVS from few quotes, eliminating daily recalibration. We introduce the Volatility Neural Process, an attention-based model that uses a two-stage training: pre-training on SABR-generated surfaces to encode a financial prior, followed by fine-tuning on market data. On S&P 500 options (2006–2023; out-of-sample 2019–2023), our model outperforms SABR, SSVI, Gaussian Process, and an ablation trained only on real data. Relative to the ablation, the SABR-induced prior reduces RMSE by about 40% and dominates in mid- and long-maturity regions where quotes are sparse. The learned prior suppresses large errors, providing a practical, data-efficient route to stable IVS construction with a single deployable model.
@@ 1 Introduction
Constructing the implied volatility surface (IVS) is a fundamental task in quantitative finance, essential for derivative pricing and risk management. A popular class of approaches are structural models (e.g., the SABR model (Hagan et al., 2002), Heston model (Heston, 1993)) and related parametric models (Gatheral, 2004; Gatheral and Jacquier, 2014). These models typically assume that the volatility of the underlying asset follows a stochastic differential equation and are effective at capturing the general shape of the volatility smile. However, their fixed mathematical forms can be overly rigid, failing to capture more complex market patterns. Another class of approaches are data-driven methods like Gaussian Process regression (De Spiegeleer et al., 2018; Roberts and Tegnér, 2021). These methods offer greater flexibility, but are prone to overfitting sparse and noisy market data, leading to financially implausible surfaces (e.g., admitting arbitrage opportunities). A common challenge of both approaches is that they treat each trading day as an isolated fitting problem, requiring daily recalibration.
@@
This limitation motivates reframing the problem as a meta-learning task: by “meta-learning” we mean training across day-level tasks to learn a procedure that reconstructs a full surface from few quotes on a new day, thereby avoiding per-day recalibration. We propose the Volatility Neural Process (VolNP), a Neural Process model (Garnelo et al., 2018b, a) designed to learn a general mapping from a sparse set of option quotes to a complete IVS. However, like other data-driven methods, Neural Processes are still prone to generating unstable and highly irregular surfaces in sparse market data. A growing body of work seeks to address this issue by incorporating financial knowledge into machine learning models, often through the imposition of constraints (e.g., no-butterfly arbitrage) derived from financial theory (Ackerer et al., 2020; Chataigner et al., 2021; Zheng et al., 2021; Hoshisashi et al., 2023; Wiedemann et al., 2025; Vuletić and Cont, 2024).
In contrast to these constraint-based methods, our approach integrates transfer learning (Chen et al., 2023). The model is first pre-trained on surfaces generated from structural models. This stage instills a prior for the shape of a financially plausible implied volatility surface into the model. Then the model is fine-tuned on real market data, enabling it to adapt this prior to complex market dynamics. This two-stage training enhances VolNP’s ability to reduce prediction errors and produce consistent surfaces from sparse data.
@@ 2 Problem Formulation
On any given trading day $q$, the implied volatility surface (IVS) can be represented by a function $f_{q}:\mathbb{R}^{2}\to\mathbb{R}^{+}$. This function maps a coordinate $\mathbf{x}=[k,\tau]^{\intercal}$ (log-moneyness and time-to-expiry) to its corresponding implied volatility $y$. We define log-moneyness as $k=\log(K/F_{q})$, where $F_{q}$ is the day $q$ forward price. The function $f_{q}$ is unobservable because only a finite set of quotes is available. Traditional methods address this by fitting a new model to these quotes each day, treating every $f_{q}$ as an independent regression problem.
We formulate the IVS construction problem within a meta-learning framework. The goal of meta-learning is to train a single, universal model that learns the process of surface construction itself. Let $\mathcal{Q}=\{1,2,\dots,Q\}$ be the set of available trading days. This model is trained on a distribution of learning tasks, where each task corresponds to constructing the IVS for a specific day $q\in\mathcal{Q}$. By learning from the wide variety of market conditions observed across $\mathcal{Q}$, the model is expected to construct a reliable implied volatility surface for a new day using only a small number of observed quotes.
A specific learning task for the model involves being given a sparse context set $\mathcal{D}_{C}=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{N}$, sampled from the market data of a single day $q\in\mathcal{Q}$, and being asked to reconstruct the entire surface $f_{q}$. Formally, the objective is to learn a single model that can use a context set $\mathcal{D}_{C}$ to produce a predictive distribution $p(\tilde{y}|\tilde{\mathbf{x}},\mathcal{D}_{C})$ for any target coordinate $\tilde{\mathbf{x}}$.
3 Volatility Neural Process
@@ 3.1 Model Architecture
To address the set-to-function regression task defined within our meta-learning framework, we propose the Volatility Neural Process (VolNP) model which is a variant of the Attentive Neural Process (Kim et al., 2019). The following describes how the model processes the inputs for a single task. Such a task consists of a context set $\mathcal{D}_{C}$ and a set of target coordinates $\mathcal{X}_{T}$. For notational simplicity, we will use these symbols throughout this subsection to refer to the inputs of a single forward pass of the model.
Figure 1: The architecture of the Volatility Neural Process (VolNP) model.
The model is an encoder-decoder architecture (see Figure 1) with attention mechanisms, designed to effectively model the complex interactions within a set of context points and reason about target predictions.
@@ 3.1.1 Encoder
The encoder’s goal is to learn the complex relationships between the individual option quotes in the context set $\mathcal{D}_{C}$. It maps this set to a matrix of contextualized representations, $\mathbf{H}^{(L)}\in\mathbb{R}^{N\times d_{r}}$, where $d_{r}$ is the hidden dimension, and each row contains a summary of a single point informed by all other points.
@@
Next, these initial representations are processed by a stack of $L$ self-attention blocks (SA). This is the core mechanism that allows the model to reason about the global structure of the volatility surface. A single option quote is only meaningful in the context of others; for instance, the implied volatility of an at-the-money option is intrinsically linked to its out-of-the-money counterparts, collectively forming the "smile" for a given maturity. The self-attention mechanism formalizes this intuition. It allows each point $\mathbf{h}_{i}^{(0)}$ to query all other points on the surface and update its representation based on their relationships. Through this process, the model learns to identify and encode complex, non-local patterns directly from the sparse data.
@@ 3.1.2 Decoder
The CA block uses the MHA, where queries come from the target representations $\mathbf{Z}$ and keys/values come from the context representation $\mathbf{H}^{(L)}$. The final target representations, $\mathbf{Z}^{(L^{\prime})}=\left[\mathbf{z}_{1}^{(L^{\prime})},\dots,\mathbf{z}_{M}^{(L^{\prime})}\right]^{\intercal}$, which contains all necessary information distilled from the context set for each specific target, are passed through a MLP, $\phi_{d}:\mathbb{R}^{d_{r}}\to\mathbb{R}^{2}$:
@@ 3.2 Training
Our model is trained using a meta-learning approach that comprises two distinct stages. The training process relies on two primary data sources.
3.2.1 Data Sources and Task Generation
The first data source is the collection of real market data, $\{\mathcal{D}_{q}\}_{q\in\mathcal{Q}}$, where $\mathcal{D}_{q}=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{N_{q}}$ is the set of all $N_{q}$ observed option quotes for a given trading day $q$. The second is a synthetic dataset, $\{\hat{\mathcal{D}}_{q}\}_{q\in\mathcal{Q}}$, where $\hat{\mathcal{D}}_{q}$ is the corresponding dense, smooth volatility surface generated by a calibrated SABR model  [^1] SABR is calibrated to each maturity slice in $\mathcal{D}_{q}$, and parameters are linearly interpolated across the term structure..
At each training step, a new task is generated. This process begins by randomly sampling a day index $q$ from the set of available trading days $\mathcal{Q}$. Subsequently, a context set $\mathcal{D}_{C}$ and a target set $\mathcal{D}_{T}$ are constructed by drawing samples from the data sources $\mathcal{D}_{q}$ and/or $\hat{\mathcal{D}}_{q}$, with the specific sampling strategy depending on the training stage.
@@ 3.2.2 Objective Function
$\mathcal{L}(\theta;\mathcal{D}_{C},\mathcal{D}_{T})=-\sum_{j=1}^{M}\log p_{\theta}(\tilde{y}_{j}|\mathcal{D}_{C},\tilde{\mathbf{x}}_{j})=\frac{1}{2}\sum_{j=1}^{M}\left(\exp(-\log\sigma_{j}^{2})(\tilde{y}_{j}-\mu_{j})^{2}+\log\sigma_{j}^{2}\right)$ (10)
@@
3.2.3 Staged Meta–Learning
The training curriculum is a two-stage process designed to first instill a robust prior and then adapt it to market specifics.
  • Stage 1: Pre-training. The goal of this stage is to instill a financial-theoretic prior into the model. For each task, a day $q$ is sampled, and the context set $\mathcal{D}_{C}$ is drawn from the sparse, real market quotes $\mathcal{D}_{q}$, while the target set $\mathcal{D}_{T}$ is drawn from the corresponding dense, synthetic SABR surface $\hat{\mathcal{D}}_{q}$. The model’s parameters $\theta$ are then updated by minimizing the loss $\mathcal{L}(\theta;\mathcal{D}_{C},\mathcal{D}_{T})$. This process encourages the model to learn a general mapping from partial, noisy observations to a globally coherent surface structure.
  • Stage 2: Fine-tuning. This stage adapts the learned prior to real-world market dynamics. The model is initialized with the parameters $\theta=(\theta_{\text{enc}},\theta_{\text{dec}})$ from the pre-training stage. For each task, a day $q$ is sampled, and now both the context set $\mathcal{D}_{C}$ and the target set $\mathcal{D}_{T}$ are sampled exclusively from the real market data $\mathcal{D}_{q}$, ensuring the sets are non-overlapping. The optimization objective remains the loss $\mathcal{L}(\theta;\mathcal{D}_{C},\mathcal{D}_{T})$.
@@ 4.1 Experimental Setup
Our experiments are carried out on a comprehensive dataset of S&P 500 (SPX) index options, spanning from January 2006 to August 2023. The raw data is from OptionMetrics dataset and is strictly partitioned by time to ensure a fair evaluation of model generalization. Data from 2006 to 2018 is used for training and validation, with 15% of the trading days randomly held out for validation. Data from 2019 to 2023 serves as the out-of-sample test set for all final evaluations. Quotes are preprocessed by computing mid-prices from best bid/ask, filter for basic liquidity. Forward prices are computed from the zero-coupon curve and dividend yields (OptionMetrics), and implied volatilities are obtained under the Black framework. Descriptive statistics of the dataset are in Table 1.
Table 1: Descriptive statistics of the dataset. The dataset is partitioned by time into a training & validation set and a test set.
Statistic Category Training & Validation Set Test Set
@@
Moneyness ATM ($|\log(K/F)|\leq 0.05$) 903,275 (50.0%) 1,183,639 (48.4%)
NTM ($0.05<|\log(K/F)|\leq 0.2$) 747,269 (41.4%) 1,047,066 (42.8%)
FTM ($|\log(K/F)|>0.2$) 156,337 (8.7%) 215,487 (8.8%)
We evaluate our proposed VolNP-FT 222‘FT’ stands for ‘Fine-tuning’. The hidden dimension of VolNP models $d_{r}$ is 128. Both the encoder and decoder MLPs ($\phi_{e},\phi_{d}$) consist of 3 hidden layers with 128 units each and GELU activation. We use $L=3$ self-attention blocks and $L^{\prime}=3$ cross-attention blocks, each with 4 attention heads. against an ablation version, VolNP-Base [^3] All models are trained with the AdamW optimizer. VolNP-FT is pre-trained for up to 200 epochs with a learning rate of $5\times 10^{-5}$ and fine-tuned for up to 100 epochs with a learning rate of $1\times 10^{-6}$. The VolNP-Base model is trained for up to 300 epochs with a learning rate of $5\times 10^{-5}$. (trained only on real data), and three benchmarks. The benchmarks are: (1) a slice-calibrated SABR model [^4] SABR is calibrated to each maturity slice, and its parameters are linearly interpolated across the term structure., (2) a Gaussian Process (GP) with an RBF kernel, and (3) a Surface SVI (SSVI) model [^5] The SSVI model uses the arbitrage-free parameterization of Gatheral and Jacquier (2014). Its global parameters are calibrated to the entire context set for each day to ensure a surface free of static arbitrage.. For each test day, we uniformly sample a context set of $N=100$ quotes from all available quotes of that day; the target set comprises the remaining quotes for that day. All baselines and VolNP variants are evaluated on the identical target sets. Performance is measured by Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE). All error metrics are reported in basis points (BPS), where 1 BPS equals 0.01% in implied volatility.

### v3

@@ Abstract
We treat implied volatility surface (IVS) reconstruction as a learning problem guided by two principles. First, we adopt a meta-learning view that trains across trading days to learn a procedure that maps sparse option quotes to a full IVS via conditional prediction, avoiding per-day calibration at test time. Second, we impose a structural prior via transfer learning: pre-train on SABR-generated dataset to encode geometric prior, then fine-tune on historical market dataset to align with empirical patterns. We implement both principles in a single attention-based Neural Process (Volatility Neural Process, VolNP) that produces a complete IVS from a sparse context set in one forward pass. On SPX options, the VolNP outperforms SABR, SSVI, and Gaussian process. Relative to an ablation trained only on market data, the SABR-induced prior reduces RMSE by about 40% and suppresses large errors, with pronounced gains at long maturities where quotes are sparse. The resulting model is fast (single pass), stable (no daily recalibration), and practical for deployment at scale.
@@ 1 Introduction
Constructing the implied volatility surface (IVS) is a fundamental task in quantitative finance, essential for derivative pricing and risk management. A popular class of approaches comprises structural models (e.g., the SABR model (Hagan et al., 2002), Heston model (Heston, 1993), Variance Gamma (Madan et al., 1998), Kou model (Kou, 2002) and rough volatility models (Bayer et al., 2016; Gatheral et al., 2022)) and related parametric models (Gatheral, 2004; Gatheral and Jacquier, 2014). These models typically assume that the underlying asset follows a stochastic differential equation and are effective at capturing the general shape of the volatility smile. However, their fixed mathematical forms can be overly rigid, failing to capture more complex market patterns. Data-driven regressors such as Gaussian process regression (De Spiegeleer et al., 2018; Roberts and Tegnér, 2021) offer flexibility but can overfit sparse and noisy quotes, producing surfaces that violate basic financial plausibility (e.g., no-arbitrage). A common limitation across both methods is to treat each trading day as an isolated fitting problem, which leads to per-day recalibration.
@@
We therefore frame IVS construction as a meta-learning problem: by meta-learning we mean training across days to learn a single procedure for few-shot conditional prediction. Concretely, we propose the Volatility Neural Process (VolNP), a neural process model (Garnelo et al., 2018a, b) that maps a small context set of option quotes to a full surface via conditional prediction. At test time, parameters are fixed (i.e., with no per-day updates), hence daily recalibration is avoided. The IVS is produced by a single forward pass given the sparse context set. However, like other data-driven methods, Neural Process is still prone to generating unstable and highly irregular surfaces in sparse market data. A growing body of work seeks to address this issue by incorporating financial knowledge into various data-driven methods, often through the imposition of constraints (e.g., no-butterfly arbitrage) derived from financial theory (Ackerer et al., 2020; Chataigner et al., 2021; Zheng et al., 2021; Hoshisashi et al., 2023; Cont and Vuletić, 2023; Ning et al., 2023; Vuletić and Cont, 2024; Wiedemann et al., 2025).
In contrast to these constraint-based methods, our approach integrates transfer learning (Chen et al., 2023). VolNP is first pre-trained on IVS generated by SABR models to encode a geometric prior for plausible volatility smiles, and is then fine-tuned on historical market data to align the prior with empirical patterns. This two-stage training (cf. Figure 1) reduces large errors, mitigates irregularities that arise in flexible models under sparsity, and lowers the incidence of butterfly-arbitrage violations.
Contributions
Our contributions are in two aspects. First, we implement a meta-learning framework to learn a procedure that can reconstruct an implied volatility surface from a limited set of quotes, eliminating the need for recalibration. For trading desks that need to update the IVS intraday, a single forward pass enables frequent surface updates, thus removing the operational bottleneck. Second, we apply transfer learning to incorporate a structural prior from SABR pre-training, followed by market fine-tuning. This reduces large errors in sparse data scenarios, improving model accuracy and robustness. VolNP processes sparse quotes and outputs a dense surface with quantified uncertainty, making it easy to integrate into existing pricing and risk management systems.
Figure 1: Flowchart for VolNP training and testing. Solid arrows show computation: historical data calibrate SABR, SABR generates synthetic IV surfaces, VolNP is pre-trained then fine-tuned, and at test time a single forward pass returns a full IV surface (no per-day calibration). Dashed arrows indicate “use as dataset” relations. In pre-training, real quotes are context and synthetic surfaces are targets; in fine-tuning, real quotes provide both context and targets.
@@ 2 Problem Formulation
On any given trading day $q$, the implied volatility surface (IVS) can be represented by a function $f_{q}:\mathbb{R}^{2}\to\mathbb{R}^{+}$. This function maps a coordinate $\mathbf{x}=[k,\tau]^{\intercal}$ (log-moneyness and time-to-expiry) to its corresponding implied volatility $y$. We define log-moneyness as $k=\log(K/F_{q})$, where $K$ is the strike price, and $F_{q}$ is the day $q$ forward price. The function $f_{q}$ is unobservable because only a finite set of market quotes is available. Let $\mathcal{Q}$ denote the index set of historical trading days used for training. Our goal is to train a single model across these days that, given a small set of quotes from day $q$, returns the conditional distribution of $y$ at any $\mathbf{x}$ on that day. This frames IVS construction as learning a procedure rather than fitting a new model every day.
For day $q\in\mathcal{Q}$, let $\mathcal{D}_{q}=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{N_{q}}$ be the finite set of $N_{q}$ observed option quotes. During training, we form two disjoint subsets from $\mathcal{D}_{q}$. The context set $\mathcal{D}_{C}\subset\mathcal{D}_{q}$ is the information that the model is allowed to condition on for that day. The target set $\mathcal{D}_{T}\subset\mathcal{D}_{q}$ is used only to define the loss and is not given to the model when making prediction. The learning objective is to train a model $p_{\theta}(y\mid\mathbf{x},\mathcal{D}_{C})$ that maps the day’s sparse context quotes to a predictive distribution at any coordinate $\mathbf{x}$, where $\theta$ denotes parameters in the model. This frames IVS construction as a set-to-function regression problem: the input is a set $\mathcal{D}_{C}$ and the output is a function on $(k,\tau)$.
In this paper, we introduce the Volatility Neural Process (VolNP) to model the conditional distribution $p_{\theta}(y\mid\mathbf{x},\mathcal{D}_{C})$. By training across many days in $\mathcal{Q}$, the VolNP learns how to use a few quotes to reconstruct a full surface. In deployment on a new day $q^{\prime}\notin\mathcal{Q}$, we often need to construct an IVS with sparse intraday quotes. The workflow is direct: use the currently available quotes as the context set $\mathcal{D}_{C}$, and evaluate the conditional distribution on any grid of $\mathbf{x}$. The model parameters remain fixed during deployment; there is no daily calibration or gradient update. A single forward pass produces a complete IVS (and associated uncertainty) for option pricing and risk management.
3 Details of Volatility Neural Process
@@ 3.1 Model Architecture
To address the set-to-function regression task defined within our meta-learning framework, we propose the Volatility Neural Process (VolNP), a variant of the Attentive Neural Process (Kim et al., 2019). We implement meta-learning and transfer learning as follows: meta-learning learns a procedure across days, while transfer learning injects a structural prior via SABR pre-training and adapts it via market fine-tuning. The following describes how the model processes the inputs for a single task. Such a task consists of a context set $\mathcal{D}_{C}$ and a set of target coordinates $\mathcal{X}_{T}$. For notational simplicity, we will use these symbols throughout this subsection to refer to the inputs of a single forward pass of the model.
Figure 2: The architecture of the Volatility Neural Process.
The model is an encoder-decoder architecture (cf. Figure 2) with attention mechanisms, designed to effectively model the complex interactions within a set of context points and reason about target predictions.
@@ 3.1.1 Encoder
The encoder’s goal is to learn the complex relationships between the individual option quotes in the context set $\mathcal{D}_{C}$. It maps this set to a matrix of contextualized representations, $\mathbf{H}^{(L)}\in\mathbb{R}^{N\times d_{r}}$, where $d_{r}$ is the hidden dimension, and each row contains a summary of a single point informed by all other points through the self-attention mechanism. Here, $N$ denotes the number of data points in $\mathcal{D}_{C}$, and $L$ indicates the number of self-attention blocks.
@@
Next, these initial representations are processed by a stack of $L$ self-attention blocks (SA). This is the core mechanism that allows the model to reason about the global structure of the volatility surface. A single option quote is only meaningful in the context of others. For instance, the implied volatility of an at-the-money option is intrinsically linked to its out-of-the-money counterparts, collectively forming the "smile" for a given maturity. The self-attention mechanism formalizes this intuition. It allows each point $\mathbf{h}_{i}^{(0)}$ to query all other points on the surface and update its representation based on their relationships. Through this process, the model learns to identify and encode complex, non-local patterns directly from the sparse data.
@@ 3.1.2 Decoder
The CA block uses the MHA, where queries come from the target representations $\mathbf{Z}$ and the Keys/Values come from the context representation $\mathbf{H}^{(L)}$. The final target representations, $\mathbf{Z}^{(L^{\prime})}=\left[\mathbf{z}_{1}^{(L^{\prime})},\dots,\mathbf{z}_{M}^{(L^{\prime})}\right]^{\intercal}$, which contains all necessary information distilled from the context set for each specific target, are passed through a MLP, $\phi_{d}:\mathbb{R}^{d_{r}}\to\mathbb{R}^{2}$:
@@ 3.2 Training
Our model is trained using an approach comprising two stages: pre-training provides a geometric prior; fine-tuning adapts this prior to real market data. The training process relies on two datasets.
3.2.1 Datasets and Task Generation
The first dataset is the collection of historical market data, $\{\mathcal{D}_{q}\}_{q\in\mathcal{Q}}$, where $\mathcal{D}_{q}=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{N_{q}}$ is the set of all $N_{q}$ observed option quotes for a given trading day $q$. The second is a synthetic dataset, $\{\hat{\mathcal{D}}_{q}\}_{q\in\mathcal{Q}}$, where $\hat{\mathcal{D}}_{q}$ is the corresponding dense, smooth volatility surface generated by a calibrated SABR model. The SABR model is calibrated to each maturity slice in $\mathcal{D}_{q}$, and parameters are linearly interpolated across the term structure.
At each training step, a new task is generated. This process begins by randomly sampling a day index $q$ from the set of available historical trading days $\mathcal{Q}$. Subsequently, a context set $\mathcal{D}_{C}$ and a target set $\mathcal{D}_{T}$ are constructed by drawing samples from the data sources $\mathcal{D}_{q}$ or $\hat{\mathcal{D}}_{q}$, with the specific sampling strategy depending on the training stage.
@@ 3.2.2 Objective Function
$\mathcal{L}(\theta;\mathcal{D}_{C},\mathcal{D}_{T})$ $=-\sum_{j=1}^{M}\log p_{\theta}(\tilde{y}_{j}\mid\mathcal{D}_{C},\tilde{\mathbf{x}}_{j})$
$=\frac{1}{2}\sum_{j=1}^{M}\left(\frac{(\tilde{y}_{j}-\mu_{j})^{2}}{\sigma_{j}^{2}}+\log\sigma_{j}^{2}\right).$
@@
3.2.3 Two-stage Training
The training curriculum is a two-stage process designed to first instill a robust prior and then adapt it to market specifics (cf. Figure 1).
  • Stage 1: Pre-training. The goal is to instill a structural, financial-theoretic prior. For each task, a day $q$ is sampled, and the context set $\mathcal{D}_{C}$ is drawn from the sparse, real market quotes $\mathcal{D}_{q}$, while the target set $\hat{\mathcal{D}}_{T}$ is drawn from the corresponding dense, synthetic SABR surface $\hat{\mathcal{D}}_{q}$. The model’s parameters $\theta$ are then updated by minimizing the loss $\mathcal{L}(\theta;\mathcal{D}_{C},\hat{\mathcal{D}}_{T})$, encoding geometric plausibility of volatility smiles.
  • Stage 2: Fine-tuning. This stage adapts the prior to real-world market dynamics. The model is initialized with the pre-trained parameters and, for each task, both the context set $\mathcal{D}_{C}$ and the target set $\mathcal{D}_{T}$ are sampled exclusively from the real market data $\mathcal{D}_{q}$, ensuring the sets are non-overlapping. The optimization objective remains $\mathcal{L}(\theta;\mathcal{D}_{C},\mathcal{D}_{T})$.
@@ 4.1 Experimental Setup
Our experiments are carried out on a dataset of S&P 500 (SPX) index options, spanning from January 2006 to August 2023. The raw data is from OptionMetrics dataset and is strictly partitioned by time to ensure a fair evaluation. Data from 2006 to 2018 is used for training and validation, with 10% of the trading days randomly held out for validation. Data from 2019 to 2023 serves as the out-of-sample test set for all final evaluations. Quotes are preprocessed by computing mid-prices from best bid/ask and applying liquidity filters: positive bid and ask with bid<<ask; daily volume >5>5; relative bid-ask spread $\leq 50\%$; and mid-price $\geq 0.1$. Forward prices are computed from the zero-coupon curve and dividend yields (from OptionMetrics), and implied volatilities are obtained under the Black framework. Descriptive statistics of the dataset are in Table 1.
Table 1: Descriptive statistics of the dataset. The dataset is partitioned by time into a training set and a test set.
Statistic Category Training Set Test Set
@@
Moneyness $|k|\leq 0.05$ 903,275 (50.0%) 1,183,639 (48.4%)
$0.05<|k|\leq 0.2$ 747,269 (41.4%) 1,047,066 (42.8%)
$|k|>0.2$ 156,337 (8.7%) 215,487 (8.8%)
We evaluate our proposed VolNP-SP 111‘SP’ stands for ‘SABR-induced Prior’. The hidden dimension of VolNP models $d_{r}$ is 128. Both the encoder and decoder MLPs ($\phi_{e},\phi_{d}$) consist of 3 hidden layers with 128 units each and GELU activation. We use $L=3$ self-attention blocks and $L^{\prime}=3$ cross-attention blocks, each with 4 attention heads. against an ablation version, VolNP-Base [^2] All VolNP models are trained with the AdamW optimizer. VolNP-SP is pre-trained for up to 200 epochs with a learning rate of $5\times 10^{-5}$ and fine-tuned for up to 200 epochs with a learning rate of $2\times 10^{-6}$. The VolNP-Base model is trained for up to 400 epochs with a learning rate of $5\times 10^{-5}$. (trained only on market data), and three benchmarks. The benchmarks are: (1) a slice-calibrated SABR model [^3] SABR is calibrated to each maturity slice, and its parameters are linearly interpolated across the term structure., (2) a Gaussian process (GP) with an RBF kernel, and (3) a Surface SVI (SSVI) model [^4] The SSVI model uses the arbitrage-free parameterization of Gatheral and Jacquier (2014). . For each test day, we uniformly sample a context set of $N=100$ quotes from all available quotes of that day; the target set comprises the remaining quotes for that day. All baselines and VolNP variants are evaluated on the identical target sets. Performance is measured by Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and coefficient of determination (R2). Let $(y_{j},\hat{y}_{j})_{j=1}^{M}$ denote true and predicted implied volatilities and let $\bar{y}=\tfrac{1}{M}\sum_{j=1}^{M}y_{j}$,
$\mathrm{MAE}=$ $\frac{1}{M}\sum_{j=1}^{M}\big|y_{j}-\hat{y}_{j}\big|$
$\mathrm{RMSE}=$ $\sqrt{\frac{1}{M}\sum_{j=1}^{M}\big(y_{j}-\hat{y}_{j}\big)^{2}}$
$R^{2}=$ $1-\frac{\sum_{j=1}^{M}\big(y_{j}-\hat{y}_{j}\big)^{2}}{\sum_{j=1}^{M}\big(y_{j}-\bar{y}\big)^{2}}$
RMSE and MAE are reported in basis points (BPS), where 1 BPS equals 0.01% in implied volatility.